OpenStudy (anonymous):

translate the sentence into a equation, and solve it. if 2 times a number is decreased by 15, the result is -3 find the number

OpenStudy (anonymous):

2x-15=-3

OpenStudy (anonymous):

2x-15=-3 To find the number: 2x=12 x=6

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

2x - 15= -3 2x= 12 x= 6